# ODE: Open-Set Evaluation of Hallucinations in Multimodal Large Language Models

### v3

@@ Abstract
Hallucination poses a persistent challenge for multimodal large language models (MLLMs). However, existing benchmarks for evaluating hallucinations are generally static, which may overlook the potential risk of data contamination. To address this issue, we propose ODE, an open-set, dynamic protocol designed to evaluate object hallucinations in MLLMs at both the existence and attribute levels. ODE employs a graph-based structure to represent real-world object concepts, their attributes, and the distributional associations between them. This structure facilitates the extraction of concept combinations based on diverse distributional criteria, generating varied samples for structured queries that evaluate hallucinations in both generative and discriminative tasks. Through the generation of new samples, dynamic concept combinations, and varied distribution frequencies, ODE mitigates the risk of data contamination and broadens the scope of evaluation. This protocol is applicable to both general and specialized scenarios, including those with limited data. Experimental results demonstrate the effectiveness of our protocol, revealing that MLLMs exhibit higher hallucination rates when evaluated with ODE-generated samples, which indicates potential data contamination. Furthermore, these generated samples aid in analyzing hallucination patterns and fine-tuning models, offering an effective approach to mitigating hallucinations in MLLMs.
Index Terms:
multimodal large language models, hallucination, evaluation protocol, open-set, dynamic.
I Introduction
Multimodal Large Language Models (MLLMs) [1, 2, 3, 4, 5, 6] have rapidly advanced in recent times, enabling detailed image descriptions (i.e., image captioning) and responses to image-related queries (i.e., visual question answering). However, these models face the persistent challenge of “hallucination”[7, 8], where generated responses appear plausible but lack fidelity to the actual image content. This issue can lead to harmful consequences, limiting the utility of MLLMs.
@@ I Introduction
Therefore, the evaluation of hallucinations in MLLMs is is crucial to improve model reliability and practical application. Prior studies have proposed various benchmarks to evaluate hallucinations in MLLMs, focusing on different types (e.g., existence hallucinations[9, 10] and relational hallucinations[11]) or levels of difficulty[12, 13, 14]. However, these benchmarks are predominantly static, using fixed test data with limited distributions, which increases the risk of data contamination. Contamination occurs when test data overlaps with training data, leading to inflated performance metrics. For example, we find that, under the same semantic distribution, models perform better on the COCO2014 image subset than on the latest Internet images (as shown in Fig. 1), with the latest internet data being less likely to be contaminated by training overlap, thus providing a more reliable evaluation. This raises concerns about whether correct responses on COCO reflect genuine understanding or result from data contamination.
@@
We argue that an effective evaluation benchmark should be open-set, meaning that evaluation data is novel to the model at both sample and distribution levels. Inspired by contamination studies in the LLM domain and guided by insights into the unique challenges of multimodal models, we outline three distinctive features that characterize our approach: (1) Out-of-distribution (OOD) evaluation at a broad distributional level rather than modifying existing datasets; (2) dynamic sample generation across different modalities; and (3) a multi-tiered dynamic structure extending from concept-level to attribute-level and distribution-level granularity.
Building on these characteristics, we introduce the Open-Set Dynamic Evaluation (ODE) protocol. ODE automatically generates datasets to evaluate object hallucinations in MLLMs, covering both existence-level and attribute-level hallucinations. Initially, we model real-world object concepts, object attributes, and concept-attribute combinations as a graph. Then, from this graph, we extract concept nodes and their associated attributes, designing diverse semantic scenes and quires for each test sample and synthesizing high-quality images. To guide the selection of concept node pairs, we design four frequency-based criteria in the following order: Standard, Long-tail, Random, and Fictional. Each criterion reflects a unique distribution pattern of object combination frequencies (see Section 2.2 for detailed explanation).
@@
  • The generated evaluation data aids model debugging, and targeted fine-tuning on ODE-identified errors or general ODE-generated samples effectively enhances model performance.
II Related Work
II-A Hallucination Evaluation in MLLMs
@@
To evaluate the degree of hallucination in MLLM, various hallucination benchmarks have been proposed. The earliest proposed CHAIR[19] measures the accuracy of object references in captions by calculating the precision of hallucinatory objects. POPE[9] improves CHIAR by evaluating the presence of hallucinations based on object detection, suitable for discriminative tasks. AMBER[10] evaluates object illusions in both discriminative and generative tasks from three dimensions: existence, attributes, and relationships. Expanding beyond object hallucinations, examples include HallusionBench [13], which focuses on visual common sense and reasoning; Hal-Eval [20], which examines event hallucinations; and CorrelationQA [21], which studies the impact of false visual inputs and other cues like counting and OCR. Some studies further evaluate complex hallucinations. VHTest[22] compares embedding similarity to find potential VH instances, while HaloQuest[14] uses real and synthesized generated images to focus on more complex scenes. Despite the proliferation of benchmarks, from simple to complex tasks, and small to large scales, testing within closed sets remains unsolved. We dynamically update the dataset by combining target concepts to achieve an open set of samples and distributions.
II-B Data Contamination
@@
Data contamination has attracted considerable attention in LLMs. The GPT-4[23] and LLama[24] reports highlight this phenomenon. Zhou et al. [15] discussed the risks and impacts of benchmark data contamination in evaluating a Master of Laws degree; similarly, Ni et al. [25] explored related concerns. Several researchers have developed methods for detecting data contamination. Fan[26], Lei[27], Zhu[17], and others introduced dynamic evaluation strategies through different algorithms to reduce data pollution. The dataset generated through the ODE protocol is dynamically distributed and can be continuously updated, which avoids the problem of data pollution that may exist in existing static datasets for illusion evaluation.
II-C Dynamic Evaluation
@@
Dynamic Evaluation refers to dynamically generating evaluation data for testing. The key to early works such as DynaBoard[28], DynaTask[29], and dynabench[30] was to utilize the intelligence of the crowd to challenge the design of evaluation sets, with the main effort focused on crowdsourcing systems and interfaces rather than algorithms, which required significant costs. Recent methods automate dynamic evaluation by rewriting and expanding samples based on task scenarios or datasets. Zhu et al.[17] used mathematical reasoning tasks to reduce data pollution in LLM by dynamically generating test samples from directed acyclic graphs. MsTemp[31], DyVal2[32], and Benchmark self evolving[33] dynamically reconstruct existing samples based on existing datasets. DeNEVIL[34] iteratively updates and improves prompts based on values, and dynamically mines samples. Our approach reduces reliance on fixed datasets by dynamically combining meta-concepts and attributes, adapting visual content and textual prompts.
III Methodology
@@
III-A Graph-Based Conceptual Modeling
Aiming to cover a broader range of target object concepts in our evaluation of object existence hallucination, ODE employs a weighted graph $G$ to model real-world scenes, facilitating the generation of more diverse scenarios in subsequent stages. Specifically, the graph $G=(V,A,E,W)$ provides an abstract representation of real-world objects, their interrelations, and attributes. Common object categories are extracted from existing datasets to form the nodes $V$, termed meta-concepts. In subsequent processes, a subset of these nodes will be selected as target concepts to test hallucination cases in the model. The attribute nodes $A$ of the main nodes $V$ represent the properties of each object concept, including state and action attributes. Additionally, quantity attributes are considered during testing. The edge weights $W$ denote the strength of the relationships between nodes, based on the co-occurrence quantity of object concepts in real-world scenes. We consider that co-occurrence frequency can reflect the semantic association and distributional characteristics between the conceptual entities $V_{i}$ and $V_{j}$. For example, the high co-occurrence frequency between ”table” and ”chair” reflects their common pairing in indoor scenes. An edge $E$ is established between nodes if a connection exists (i.e., the edge weight $W$ is non-zero). To facilitate a more detailed analysis of the mutual influence of illusions among these concepts, we categorize the concepts into environment-level $V_{\text{env}}$ and entity-level $V_{\text{ent}}$, such as ”grass” and ”frisbee”. We particularly focus on two coexistence patterns: entity-environment and entity-entity, ensuring the content generated in subsequent stages targeted.
ODE performs concept extraction from real-world scenarios and then focus on hallucinatory scenarios by generating a graph that captures hallucination associations, facilitating the creation of a comprehensive range of test samples for general hallucination evaluation. This modeling approach is applicable to both general and specific domains. We further emphasize the customization of hallucination detection tailored to specific fields.
III-B Construction of Semantic Scenes
@@
III-B1 Selection Criteria
Through extensive learning from multimodal data, the model gradually acquires semantic representations of concepts; however, it may show varying comprehension of concept combinations, especially with diverse distributions, potentially causing hallucinations. Based on this observation, we design four distributional levels of concept combination standards—Standard, Long-tail, Random, and Fictional—to evaluate the model’s grasp of different association patterns.
@@ III-B1 Selection Criteria
    where $\epsilon$ and $\delta$ set co-occurrence bounds. This standard assesses the model’s performance on rare, associative pairs. Attributes are also chosen from commonly observed attributes to align with realistic contexts.
@@ III-B2 Scene Content
III-C Generation and Filtering of Images
@@
III-D Structuring of Inquiries
We develop an evaluation Inquiry template specifically designed to evalute object-level and attribute-level hallucinations (state, action, and quantity) through automated generation.
For generative tasks, we use the inquiry “Please describe this image.” to instruct the MLLM to identify the image’s concepts. For discriminative tasks, object existence hallucinations are evaluated with inquiry like “Is there a {object} in the image?” Attribute hallucinations are evaluated with inquiries like “Is the {object} {state attribute} in the image?”, ”Does the {object} {action attribute} in the picture?”, and ”Is/Are there {quantity} {object(s)} in the picture?”, expecting a “yes” or “no” response. To examine hallucinated objects, we include counterfactual prompts in the discriminative task, inquiring about nonexistent objects or attributes, such as “Is there a {hallucinated object} in the image?”
TABLE I: Evaluation results of different models on both generative and discriminative tasks across various scenarios. Highlighted items indicate optimal performance. Yellow highlights denote better performance with lower values, while blue highlights indicate better performance with higher values.
@@ III-D Structuring of Inquiries
IV Experiments
IV-A Setup
Data Preparation. We extracted real-world target concepts through the AMBER benchmark[10], a multidimensional benchmark for evaluating MLLM hallucinations that includes 337 objects across 14 domains reflecting standard concepts After concept modeling, we selected 40 concept combinations at four different difficulty distributional levels to generate images, resulting in 808 test images after sampling. Each image was paired with factual and hallucination questions, creating 8,786 test data pairs for both discriminative and generative tasks.
MLLMs Evaluated. We selected several state-of-the-art MLLMs for evaluation, including MiniGPT-4[4], InstructBLIP[3], LLaVA-1.5[2], CogVLM[6], and mPLUG_Owl[5]. To ensure fair evaluation, we used each model’s official hyperparameters to avoid length-based bias in response generation.
Evaluation Metrics. We adopted AMBER-derived evaluation metrics. For generative tasks, CHAIR measured hallucination frequency, Cover assessed content coverage, Hal represented the proportion of hallucinated responses, and Cog evaluated similarity between generated hallucinations and human cognitive hallucinations. For discriminative tasks, we used standard classification metrics: Accuracy, Precision, Recall, and F1-Score. To evaluate hallucinations, ODE calculates Precision and Recall specifically for hallucination-related questions (ground truth ”no”) and uses Accuracy across all questions to prevent models from achieving high scores by simply rejecting answers.
@@ IV-A Setup
IV-B Effectiveness of Synthetic Images
@@
IV-C Main Results
Table I shows our test results (refer to the Appendix for details), and we found that:
  • Inconsistencies Between Static Benchmarks and ODE Performance: Models like MiniGPT-4 and LLaVA perform well on static benchmarks, with F1-scores around 80-90. However, when tested with standard concept pairs in our evaluation, their scores decline. This suggests that these models may not fully understand fundamental concept relationships but rather rely on memorized correlations or noise from training data. This outcome highlights the limitations of static benchmarks, which can overestimate model capabilities due to data leakage and overfitting. Dynamic testing with diverse samples is needed for a deeper evaluation of model learning and adaptability.
@@ IV-C Main Results
  • Different Requirements of Tasks on Semantic Understanding: Our results reveal distinct adaptive differences in model performance between high-frequency and low-frequency concepts in generative vs. discriminative tasks. During the training phase, models are repeatedly exposed to these high-frequency combinations, establishing robust semantic associations. This reinforced in generative tasks which emphasize semantic consistency. This flexibility enables generative models to produce outputs that broadly align with concepts, even with incomplete understanding of details. Discriminative tasks, however, require precise identification of target concepts, making models sensitive to noise or biases from overrepresented pairs in training data. This over-memorization affects accuracy more in discriminative tasks, where even small misclassifications disrupt coherence.
  • Widening Performance Differences Among Models: Different architectures and training methods, such as memory mechanisms and attention distributions, lead to performance variations. For example, mPLUG performs well in attribute discrimination, while MiniGPT-4 maintains balanced performance across tasks but lacks fine-grained attribute precision. Overall, LLaVA-1.5 exhibits robust performance across tasks and difficulty levels, while models like MiniGPT-4 and CogVLM highlight distinct strengths and weaknesses in data dependence and generalization capabilities.
TABLE II: Performance of models on both generative and discriminative tasks (object existence hallucination) across two datasets, with the difference ($\Delta$) between datasets.
@@
IV-D Dynamic Benchmarking Updates
As a third-party benchmark evaluation, to ensure that models maintain robust performance in evolving application scenarios, we propose an ODE-based dynamic update mechanism for benchmark evaluations. This update mechanism involves the introduction of new target concept pairs, transformations in attribute combinations, and updates to the underlying knowledge base to include a broader array of object categories and their relationships. Such a iteratively update process enables the benchmark to cover a wider range of distributional scenarios, ensuring both the accuracy and comprehensiveness of model evaluation results. We conducted a second evaluation using data generated by Stable-Diffusion 1.5. Table II presents the model’s performance variations under a long-tail distribution. The significant discrepancies in performance indicate differing levels of the model’s comprehension across various concepts.
V Methodological Applications
V-A Hallucination Tendencies in Evaluation Results
Our findings enable analysis of hallucination tendencies within individual concepts and hallucination associations between different concepts. For example, we generated a frequency matrix of fact-hallucination concept pairs from LLaVA and performed clustering, identifying four groups, including indoor and traffic scene concepts. We found that hallucinations are more likely in scenarios with high contextual similarity or visual ambiguity. For instance, clusters containing both indoor and outdoor concepts (e.g., “car” and “chair”) exhibit a higher hallucination rates, as shown in Table III, revealing potential weaknesses in the model’s understanding of scene context and object differentiation.
TABLE III: Cluster Analysis: Top Truth Concepts

### v4

@@ Abstract
Hallucination poses a persistent challenge for multimodal large language models (MLLMs). However, existing benchmarks for evaluating hallucinations are generally static, which may overlook the potential risk of data contamination. To address this issue, we propose ODE, an open-set, dynamic protocol designed to evaluate object hallucinations in MLLMs at both the existence and attribute levels. ODE employs a graph-based structure to represent real-world object concepts, their attributes, and the distributional associations between them. This structure facilitates the extraction of concept combinations based on diverse distributional criteria, generating varied samples for structured queries that evaluate hallucinations in both generative and discriminative tasks. Through the generation of new samples, dynamic concept combinations, and varied distribution frequencies, ODE mitigates the risk of data contamination and broadens the scope of evaluation. This protocol is applicable to both general and specialized scenarios, including those with limited data. Experimental results demonstrate the effectiveness of our protocol, revealing that MLLMs exhibit higher hallucination rates when evaluated with ODE-generated samples, which indicates potential data contamination. Furthermore, these generated samples aid in analyzing hallucination patterns and fine-tuning models, offering an effective approach to mitigating hallucinations in MLLMs. Our code are available at https://github.com/Iridescent-y/ODE.
1 Introduction
Multimodal Large Language Models (MLLMs) [1, 16, 3, 36, 33, 30] have rapidly advanced in recent times, enabling detailed image descriptions (i.e., image captioning) and responses to image-related queries (i.e., visual question answering). However, these models face the persistent challenge of “hallucination” [15, 8, 27], where generated responses appear plausible but lack fidelity to the actual image content. This issue can lead to harmful consequences, limiting the utility of MLLMs.
@@ 1 Introduction
Therefore, the evaluation of hallucinations in MLLMs is crucial to improve model reliability and practical application. Prior studies have proposed various benchmarks to evaluate hallucinations in MLLMs, focusing on different types (e.g., existence hallucinations[14, 26] and relational hallucinations [34]) or levels of difficulty [32, 6, 31]. However, these benchmarks are predominantly static, using fixed test data with limited distributions, which increases the risk of data contamination. Contamination occurs when test data overlaps with training data, leading to inflated performance metrics. For example, we find that, under the same semantic distribution, models perform better on the COCO2014 image subset than on the latest Internet images (as shown in Fig. 1), with the latest Internet data being less likely to be contaminated by training overlap, thus providing a more reliable evaluation. This raises concerns about whether correct responses on COCO reflect genuine understanding or result from data contamination.
@@
We argue that an effective evaluation benchmark should be open-set, meaning that evaluation data are novel to the model at both sample and distribution levels. Inspired by contamination studies in the LLM domain and guided by insights into the unique challenges of multimodal models, we outline three distinctive features that characterize our approach: (1) Out-of-distribution (OOD) evaluation at a broad distributional level rather than modifying existing datasets; (2) dynamic sample generation across different modalities; and (3) a multi-tiered dynamic structure extending from concept-level to attribute-level and distribution-level granularity.
Building on these characteristics, we introduce the Open-Set Dynamic Evaluation (ODE) protocol. ODE automatically generates datasets to evaluate object hallucinations in MLLMs, covering both existence-level and attribute-level hallucinations. Initially, we model real-world object concepts, object attributes, and concept-attribute combinations as a graph. Then, from this graph, we extract concept nodes and their associated attributes, designing diverse semantic scenes and queries for each test sample and synthesizing high-quality images. To guide the selection of concept node pairs, we design four frequency-based criteria in the following order: Standard, Long-tail, Random, and Fictional. Each criterion reflects a unique distribution pattern of object combination frequencies (see Section 2.2 for detailed explanation).
@@
  • The generated evaluation data aid model debugging, and targeted fine-tuning on ODE-identified errors or general ODE-generated samples effectively enhances model performance.
2 Related Work
2.1 Hallucination Evaluation in MLLMs
@@
To evaluate the degree of hallucination in MLLMs, various hallucination benchmarks have been proposed. The earliest proposed CHAIR [22] measures the accuracy of object references in captions by calculating the precision of hallucinatory objects. POPE [14] improves CHAIR by evaluating the presence of hallucinations based on object detection, suitable for discriminative tasks. AMBER [26] evaluates object hallucinations in both discriminative and generative tasks from three dimensions: existence, attributes, and relationships. Expanding beyond object hallucinations, examples include HallusionBench [6], which focuses on visual common sense and reasoning; Hal-Eval [10], which examines event hallucinations; and CorrelationQA [7], which studies the impact of false visual inputs and other cues like counting and OCR. Some studies further evaluate complex hallucinations. VHTest [9] compares embedding similarity to find potential VH instances, while HaloQuest [31] uses real and synthesized generated images to focus on more complex scenes. Despite the proliferation of benchmarks, from simple to complex tasks, and small to large scales, testing within closed sets remains unsolved. We dynamically update the dataset by combining target concepts to achieve an open set of samples and distributions.
2.2 Data Contamination
@@
Data contamination has attracted considerable attention in LLMs. The GPT-4[19] and LLama [25] reports highlight this phenomenon. Zhou et al. [35] discussed the risks and impacts of benchmark data contamination in evaluating LLMs; similarly, Ni et al. [21] explored related concerns. Several researchers have developed methods for detecting data contamination. Fan[5], Lei [12], Zhu [37], and others introduced dynamic evaluation strategies through different algorithms to reduce data pollution. The dataset generated through the ODE protocol is dynamically distributed and can be continuously updated, which avoids the problem of data pollution that may exist in existing static datasets for illusion evaluation.
2.3 Dynamic Evaluation
@@
Dynamic Evaluation refers to dynamically generating evaluation data for testing. The key to early works such as DynaBoard [20], DynaTask [23], and Dynabench [11] was to utilize the intelligence of the crowd to challenge the design of evaluation sets, with the main effort focused on crowdsourcing systems and interfaces rather than algorithms, which required significant costs. Recent methods automate dynamic evaluation by rewriting and expanding samples based on task scenarios or datasets. Zhu et al. [37] used mathematical reasoning tasks to reduce data pollution in LLM by dynamically generating test samples from directed acyclic graphs. MsTemp [17], DyVal2 [38], and Benchmark self evolving[29] dynamically reconstruct samples from existing datasets. DeNEVIL [4] iteratively updates and improves prompts based on values, and dynamically mines samples. Our approach reduces reliance on fixed datasets by dynamically combining meta-concepts and attributes, adapting visual content and textual prompts.
3 Methodology
@@
3.1 Graph-Based Conceptual Modeling
Aiming to cover a broader range of target object concepts in our evaluation of object existence hallucination, ODE employs a weighted graph $G$ to model real-world scenes, facilitating the generation of more diverse scenarios in subsequent stages. Specifically, the graph $G=(V,A,E,W)$ provides an abstract representation of real-world objects, their interrelations, and attributes. Common object categories are extracted from existing datasets to form the nodes $V$, termed meta-concepts. In subsequent processes, a subset of these nodes will be selected as target concepts to test hallucination cases in the model. The attribute nodes $A$ of the main nodes $V$ represent the properties of each object concept, including state and action attributes. Additionally, quantity attributes are considered during testing. The edge weights $W$ denote the strength of the relationships between nodes, based on the co-occurrence quantity of object concepts in real-world scenes. We consider that co-occurrence frequency can reflect the semantic association and distributional characteristics between the conceptual entities $V_{i}$ and $V_{j}$. For example, the high co-occurrence frequency between ”table” and ”chair” reflects their common pairing in indoor scenes. An edge $E$ is established between nodes if a connection exists (i.e., the edge weight $W$ is non-zero). To facilitate a more detailed analysis of the mutual influence of hallucinations among these concepts, we categorize the concepts into environment-level $V_{\text{env}}$ and entity-level $V_{\text{ent}}$, such as ”grass” and ”frisbee”. We particularly focus on two coexistence patterns: entity-environment and entity-entity coexistence patterns, ensuring the content generated in subsequent stages targeted.
ODE performs concept extraction from real-world scenarios and then focuses on hallucinatory scenarios by generating a graph that captures associations related to hallucinations, facilitating the creation of a comprehensive range of test samples for general hallucination evaluation. This modeling approach is applicable to both general and specific domains. We further emphasize the customization of hallucination detection tailored to specific fields.
3.2 Construction of Semantic Scenes
@@
3.2.1 Selection Criteria
Through extensive learning from multimodal data, the model gradually acquires semantic representations of concepts, however, it may show varying comprehension of concept combinations, especially with diverse distributions, potentially causing hallucinations. Based on this observation, we design four distributional levels of concept combination standards—Standard, Long-tail, Random, and Fictional—to evaluate the model’s grasp of different association patterns.
@@ 3.2.1 Selection Criteria
    where $\epsilon$ and $\delta$ set co-occurrence bounds. This standard evaluates the performance of the model in rare, associative pairs. Attributes are also chosen from commonly observed attributes to align with realistic contexts.
@@ 3.2.2 Scene Content
3.3 Generation and Filtering of Images
@@
3.4 Structuring of Inquiries
We develop an evaluation Inquiry template specifically designed to evaluate object-level and attribute-level hallucinations (state, action, and quantity) through automated generation.
For generative tasks, we use the inquiry “Please describe this image.” to instruct the MLLM to identify the image’s concepts. For discriminative tasks, object existence hallucinations are evaluated with inquiries like “Is there a {object} in the image?” Attribute hallucinations are evaluated with inquiries like “Is the {object} {state attribute} in the image?”, ”Does the {object} {action attribute} in the picture?”, and ”Is/Are there {quantity} {object(s)} in the picture?”, expecting a “yes” or “no” response. To examine hallucinated objects, we include counterfactual prompts in the discriminative task, inquiring about nonexistent objects or attributes, such as “Is there a {hallucinated object} in the image?”
Table 1: Evaluation results of different models on both generative and discriminative tasks across various scenarios. Highlighted cells indicate optimal performance. Yellow highlights denote better performance with lower values, while blue highlights indicate better performance with higher values.
@@ 3.4 Structuring of Inquiries
4 Experiments
4.1 Setup
Data Preparation. We extracted real-world target concepts through the AMBER benchmark[26], a multidimensional benchmark for evaluating MLLM hallucinations that includes 337 objects across 14 domains reflecting standard concepts. After concept modeling, we selected 40 concept combinations at four different distributional difficulty levels to generate images, resulting in 808 test images after sampling. Each image was paired with factual and hallucination questions, creating 8,786 test data pairs for both discriminative and generative tasks.
MLLMs Evaluated. We selected several state-of-the-art MLLMs for evaluation, including MiniGPT-4 [36], InstructBLIP [3], LLaVA-1.5 [16], CogVLM [30], mPLUG_Owl [33], Cambrian [24], InternVL-2.5 [2], and Qwen2-VL [28]. To ensure fair evaluation, we used each model’s official hyperparameters to avoid length-based bias in response generation.
Evaluation Metrics. We adopted AMBER-derived evaluation metrics. For generative tasks, CHAIR measured hallucination frequency, Cover assessed content coverage, Hal represented the proportion of hallucinated responses, and Cog evaluated the similarity between generated hallucinations and human cognitive hallucinations. For discriminative tasks, we used standard classification metrics: Accuracy, Precision, Recall, and F1-Score. To evaluate hallucinations, ODE calculates Precision and Recall specifically for hallucination-related questions (ground truth ”no”) and uses Accuracy across all questions to prevent models from achieving high scores by simply rejecting answers.
@@ 4.1 Setup
4.2 Effectiveness of Synthetic Images
@@
4.3 Main Results
Table 1 shows selected evaluation results (complete data are provided in the Appendix), and we found that:
  • Inconsistencies Between Static Benchmarks and ODE Performance: Models like MiniGPT-4 and LLaVA perform well on static benchmarks, with F1-scores around 80-90. However, when tested with standard concept pairs in our evaluation, their scores declined. This suggests that these models may not fully understand fundamental concept relationships but rather rely on memorized correlations or noise from training data. This outcome highlights the limitations of static benchmarks, which can overestimate model capabilities due to data leakage and overfitting. Dynamic testing with diverse samples is necessary for a deeper evaluation of model learning and adaptability.
@@ 4.3 Main Results
  • Different Requirements of Tasks on Semantic Understanding: Our results reveal distinct adaptive differences in model performance between high-frequency and low-frequency concepts in generative vs. discriminative tasks. During the training phase, models are repeatedly exposed to these high-frequency combinations, establishing robust semantic associations. This is reinforced in generative tasks which emphasize semantic consistency. This flexibility enables generative models to produce outputs that broadly align with concepts, even with incomplete understanding of details. Discriminative tasks, however, require precise identification of target concepts, making models sensitive to noise or biases from overrepresented pairs in training data. This over-memorization affects accuracy more in discriminative tasks, where even small misclassifications disrupt coherence.
  • Widening Performance Differences Among Models: Different architectures and training methods, such as memory mechanisms and attention distributions, lead to performance variations. For example, mPLUG performs well in attribute discrimination, while MiniGPT-4 maintains balanced performance across tasks but lacks fine-grained attribute precision. Overall, LLaVA-1.5 exhibits robust performance across tasks and difficulty levels, while models like MiniGPT-4 and CogVLM highlight distinct strengths and weaknesses in data dependency and generalization capabilities.
Table 2: Performance of models on both generative and discriminative tasks (object existence hallucination) across two datasets, with the difference ($\Delta$) between datasets.
@@
4.4 Dynamic Benchmarking Updates
As a third-party benchmark evaluation, to ensure that models maintain robust performance in evolving application scenarios, we propose an ODE-based dynamic update mechanism for benchmark evaluations. This update mechanism involves the introduction of new pairs of target concepts, transformations in attribute combinations, and updates to the underlying knowledge base to include a broader range of object categories and their relationships. Such an iterative update process enables the benchmark to cover a wider range of distributional scenarios, ensuring both the accuracy and the comprehensiveness of model evaluation results. We conducted a second evaluation using data generated by Stable-Diffusion 1.5. Table 2 presents the model performance variations under a long-tail distribution. The significant discrepancies in performance indicate differing levels of model comprehension across various concepts.
5 Methodological Applications
5.1 Hallucination Tendencies in Evaluation Results
Our findings enable the analysis of hallucination tendencies within individual concepts and hallucination associations between different concepts. For example, we generated a frequency matrix of fact-hallucination concept pairs from LLaVA and performed clustering, identifying four groups, including indoor and traffic scene concepts. We found that hallucinations are more likely in scenarios with high contextual similarity or visual ambiguity. For instance, clusters containing both indoor and outdoor concepts (e.g., “car” and “chair”) exhibit higher hallucination rates, as shown in Table 3, revealing potential weaknesses in the model’s understanding of scene context and object differentiation.
Table 3: Cluster Analysis: Top Truth Concepts